A one sample +-test is conducted

on Ho: M = 81.6. The sample has a

mean of 84.1, s = 3.1, n = 25, and a =

01. What conclusion can be drawn?


1
Expert's answer
2022-06-16T09:49:08-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=81.6"

"H_1:\mu\not=81.6"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=24" and the critical value for a two-tailed test is "t_c =2.79694."

The rejection region for this two-tailed test is "R = \{t:|t|>2.79694\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{84.1-81.6}{3.1/\sqrt{25}}=4.0323"


Since it is observed that "|t|=4.0323>2.79694=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=4.0323" is "p= 0.000486," and since "p=0.000486<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 81.6, at the "\alpha = 0.01" significance level.